# Perception of Visual Content: Differences between Humans and Foundation Models

### v2

@@ Abstract
Human-annotated content is often used to train machine learning (ML) models. However, recently, language and multi-modal foundational models have been used to replace and scale-up human annotator’s efforts. This study compares human-generated and ML-generated annotations of images representing diverse socio-economic contexts. We aim to understand differences in perception and identify potential biases in content interpretation. Our dataset comprises images of people from various geographical regions and income levels, covering various daily activities and home environments. We compare human and ML-generated annotations semantically and evaluate their impact on predictive models. Our results show highest similarity between ML captions and human labels from a low-level perspective, i.e., types of words that appear and sentence structures, but all three annotations are alike in how similar or dissimilar they perceive images across different regions. Additionally, ML Captions resulted in best overall region classification performance, while ML Objects and ML Captions performed best overall for income regression. The varying performance of annotation sets highlights the notion that all annotations are important, and that human-generated annotations are yet to be replaceable.
@@ Introduction
In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images. This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect unbiasedness in predictive models. This leads to our second research question (RQ2): “How do different combinations of annotations affect the performance and bias in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models. These annotations are based on images of people performing different activities from varying geographical regions and socio-economic classes.
@@
Figure 1: Handwashing images annotated using the pre-trained Faster R-CNN model.
@@
The dataset utilised in our study was obtained from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021). Dollar Street visualises global public data by arranging families based on their income levels, from the poorest on the left to the wealthiest on the right. It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes. The dataset contains 3,041 images, sourced from various geographical regions and income levels, consisting of the following image categories: Drinking Water, Drying Clothes, Front Doors, Hand Washing, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes. These images include a diverse range of topics, covering both actions and static elements, such as home environments. This is to provide a good representation of different aspects of daily life, which is applicable across all cultures and income levels.
@@
Prior to obtaining all three annotation sets, the images were selected based on whether they successfully led to the generation of ML Objects with at least 50% confidence and ML Captions beforehand. Moreover, duplicates were removed and down-sampling was performed to ensure the images were balanced for all regions and income groups. The human annotations were processed to deal with inconsistencies. These steps involve noise/punctuation removal, conversion to lower case, auto-correcting misspellings, and segmenting strings that are concatenated. The number of images in the dataset decreased at each step, starting at 3,041 images at the start, reducing to 2,797 after generating the ML Objects and Captions, then to 1,888 after removing duplicates and down-sampling, and finally to 1,886 unique samples after crowdsourcing. In total, 1,155 samples were discarded from the original dataset.
We utilised a fine-tuned sentence transformer model to convert the three types of annotations into sentence embeddings. We opted to fine-tune an existing model instead of using an pre-trained model out of the box as we want the model and representations to be very focused to our data. We fine-tuned a pre-trained Sentence Transformer model, “all-MiniLM-L12-v2”, available at the Hugging Face Hub [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2. The fine-tuning process was run over 5 epochs with a batch size of 16. The model was fine-tuned on the three annotation sets so that it can better capture the characteristics of the annotations, hence generating closer embeddings. The generated embeddings have a size of 384 dimensions, which is the default embedding size from the original pre-trained model.
Exp. 1: Investigating Annotation Similarity (RQ1)
@@
Exp. 2: Exploring How Annotations Impact Performance and Bias in Predictive Models (RQ2)
This section emphasises the diversity of model prediction outcomes over overall model effectiveness. These experiments aim to evaluate the distribution of prediction errors across different regions and categories, allowing us to identify potential discrimination, e.g., higher prediction errors for certain subsets of the population as compared to the error on others. To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
@@
Figure 3 illustrates a t-SNE visualisation of the vector embeddings of the three annotation categories for each data sample. The results show how, despite having a fine-tuned embedding model, the embeddings form into two distinct clusters, with one for ML Captions and Human labels and another for ML Objects. This evidence implies that there exist differences between the different types of annotations we collected in the study. The clusters of embeddings are due to differences in the annotations. Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images. On the other hand, the ML-generated object labels are primarily object-centric. The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification. The ML-generated captions would be more descriptive than the object labels and although they may not be as specific as the human annotations, they are more likely to contain actions and objects, which may also be present in Human Labels. Additionally, there are apparent syntactic differences between the annotations - the ML Objects consists of lists of objects per image; the ML Captions consist of a phrase per image; while the Human Labels are a list of words and phrases. It is also important to note that a limitation that could have affected this categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information. This causes the sentence structures in the annotations to remain unchanged, resulting in more pronounced categorisations. Considering these differences, the clusters in the visualisation are expected.
@@
Since t-SNE is only a two-dimensional representation of the data, it is useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset. This is shown in Table 1. These values give a more accurate and quantitative comparison of the annotation similarities. It can be seen that the ML Captions and Human Labels are the most similar pair of annotations (0.69), followed by ML Objects and ML Captions (0.50), and ML Objects and Human Labels (0.42). It is important to note that when comparing ML Objects to any of the other annotations, the similarity scores are relatively low. To further investigate length statistics per image, Human Labels averaged 7.18 words per worker, ML Captions averaged 8.62 words per caption, and ML Objects averaged 4.36 object labels per prediction. The length of ML Objects is determined by the number of objects in the predictions rather than the average number of words as ML Objects consists of a discrete list of items.
@@
To further explore the relationships of these similarity scores, a plot of pairwise relationships is presented in figure 4. The plot illustrates how similarity scores between two annotation categories correlate to another pair’s scores for a given image. The plot is colour-coded by region to allow further analysis of similarity scores by geographical location. The analysis reveals that most regions show no particular trends or patterns, with one exception: images from Europe tend to cluster in the higher similarity range (top right area) of the plots, indicating stronger agreement between all annotation types for European images. Moreover, when looking at the subplots for ML Captions-Human Labels pairs, i.e., subplots on the third column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers. This aligns with the previous observation that ML Captions and Human Labels have the highest average similarity score (0.69). It can also be seen that all three pairs show the same amount of spread of scores across all subplots, which correspond to the variance values of the mean similarity scores between each pair (0.12).
@@
Moreover, the fact that the ML Objects annotations have a relatively low similarity score indicates that the ML object detection model may be detecting objects that do not appear in the corresponding ML Captions or Human Labels, or failing to identify objects that do. This reasoning is consistent with the findings of Zhang et al. (2021), which suggest that deeper neural networks, such as ResNet - the architecture of our object detection model - may struggle more with difficult images. Perhaps there is obstruction or lack of objects in the image, leading to no or incorrect objects being detected in these outliers, resulting in the low similarity scores when compared to their respective ML Captions and Human Labels. The ML captioning model as well as human annotators have the advantage in that they are able to capture non-object elements, such as actions, e.g., “wash”, or environmental descriptions, e.g., “dirty”.
@@
The variation in optimal models across the image categories is understandable. Images from each of the categories may contain varying objects and/or actions that are associated with certain geographical regions, e.g., for theHand Washing category, images of buckets may appear more frequently in samples from Africa, whereas images of sinks may appear more in samples from Europe. This explains why certain annotations are better than others at capturing relationships between the objects or actions and the regions. The best-performing models for five of the categories (Drying Clothes, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes) are the ones with a combination of at least two annotation sets. This may be due to how the images from these categories may have discriminative features that cannot be captured with just one annotation set.
Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the ML Captions Model and the Human Labels Model at $p$-value $<$ $1.15\times 10^{-7}$. This statistically significant value suggests that the ML Captions Model’s better performance is statistically reliable.
@@
It is apparent that, unlike the classification model, the regression model with ML Objects and ML Captions performed the best overall, indicated by its lowest RMSE value (1817.49). The best-performing models differ across all regions. The model containing all annotations was the most effective for African images (561.42), similar to the classification model, while the Asian images achieved the best results with the Human Labels Model (1408). Europe saw optimal performance from the ML Objects and ML Captions Model (2777.27), and the Americas was best handled by the ML Captions and Human Labels Model (1578.29). Moreover, it is worth highlighting that Asia is the only region where a stand-alone annotation set achieved the best performance.
Looking at the performance of the models by category, it is evident that the ML Objects-ML Captions Model is the most consistent and achieves the best performance across half of the categories, namely Drinking Water, Drying Clothes, Kitchens, and Washing Clothes. The ML Objects-Human Labels Model was the most effective for Front Doors. The Handwashing category was best handled by the ML Captions Model, while the Human Labels Model achieved the best results for Living Rooms and Places for Dinner.
@@
The observation that the ML Objects and ML Captions Model achieves highest performance aligns with the claim made by Zhang et al. (2021) on machines being better at selecting distinguishing features in the context of image segmentation. In our case of income regression, this also applies, although only for ML Objects rather than ML Captions. This could be attributed to certain objects in the images being associated with certain income levels, therefore improving the model’s ability to predict income. Further experiments would need to be conducted to explore this. Moreover, similar to the classification model, the all-annotations model worked best for African images, likely driven by the need for the annotations to encompass various features, including objects and actions. The images from Asia and the Americas were best handled by the Human Labels Model and the ML Captions-Human Labels Model, respectively. This shows how images from both regions have features indicative of income that are most likely not object-based, rather based on actions or nuanced descriptions of the environment. The fact that Asia required only Human Labels for its optimal performance suggests that the discriminative features for income still require full human intervention, and that although the ML annotations can still capture them, they do not perform as well, even when combined with the Human Labels. In contrast, the European images achieved best performance with the ML Objects-ML Captions Model. This suggests that the combined ML-based annotations are better at capturing the defining features and relationships to predict income for European images. This indicates a potential for bias in the ML annotations, potentially stemming from the training data of the ML models, which may be more representative of European content. However, another possibility is that the human annotators coincidentally managed to find better discriminating features for income prediction when annotating images from Asia.
@@
The best regression models for the action-based image categories are either the ML Captions Model or the ML Objects-ML Captions Model, indicating how the ML models are better at finding the defining features to predict income. It may be due to how the images contain more easily identifiable actions and objects that are indicative of income. On the other hand, image categories that are not action-based, such as Living Rooms and Places for Dinner are best handled by the Human Labels Model. This suggests that humans are better at identifying key features for income prediction when said features are likely to be based on nuanced descriptions.
@@
Figure 5 illustrates a scatter plot of the regression model trained using the ML Objects and ML Captions to predict income on the test dataset. Each data point represents an individual image sample from the dataset, which corresponds to a household. The points include associated attributes, such as region (continent), country, actual income, and predicted income. They are also colour-coded by continent. The blue dotted line represents the line of best fit and the red solid line represents the identity line. The x-axis represents the ground truth of the income, which has been plotted in log scale for a clearer visualisation, whereas the y-axis represents the predicted income, plotted in its original value.
The plot reveals that the samples whose incomes are underestimated come from all regions, with most of the Europe samples clustering towards the higher income area in the x-axis and the Africa samples clustering towards the lower income area in the x-axis. On the other hand, there are very few overestimated samples, most of which are only from Africa.
The underestimation of samples may be attributed to the high number of samples within the lower to middle income range in the dataset. Interestingly, when looking at the distribution, samples from Africa tend to be closer to the red identity line, indicating a more accurate prediction, especially compared to Europe. This is expected as this is also reflected in the RMSE scores of the regression models in Table 4. The low variance of the RMSE values for African samples may be due to the features and relationships in the African samples captured in the ML Objects and ML Captions being more different from those in other regions, enabling the model to be more consistent in its predictions. On the other hand, the key features found in the annotations of the European samples may be more complex, resulting in more variability in predictions, as shown in the vertical spread of the European samples in the scatter plot, as well as the greater variance in the RMSE table.

### v3

@@ Abstract
Human-annotated content is often used to train machine learning (ML) models. However, recently, language and multi-modal foundational models have been used to replace and scale-up human annotator’s efforts. This study explores the similarity between human-generated and ML-generated annotations of images across diverse socio-economic contexts (RQ1) and their impact on ML model performance and bias (RQ2). We aim to understand differences in perception and identify potential biases in content interpretation. Our dataset comprises images of people from various geographical regions and income levels, covering various daily activities and home environments. ML captions and human labels show highest similarity at a low-level, i.e., types of words that appear and sentence structures, but all annotations are consistent in how they perceive images across regions. ML Captions resulted in best overall region classification performance, while ML Objects and ML Captions performed best overall for income regression. ML annotations worked best for action categories, while human input was more effective for non-action categories. These findings highlight the notion that both human and machine annotations are important, and that human-generated annotations are yet to be replaceable.
Code and Datasetshttps://github.com/nardienapratama/Human-Machine-Perception-of-Visual-Content
@@ Introduction
In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images. This leads to our first research question (RQ1): “How similar are human-generated and ML-generated annotations?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect unbiasedness in predictive models. This leads to our second research question (RQ2): “How do different combinations of annotations affect the performance and bias in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models. These annotations are based on images of people performing different activities from varying geographical regions and socio-economic classes.
@@
Figure 1: Example images annotated with the pre-trained Faster R-CNN model. Left: Hand Washing. Right: Living Rooms.
@@
The dataset utilised in this study comes from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021). Dollar Street visualises global public data by arranging families based on their income levels. It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes. The dataset contains 3,041 images, sourced from various geographical regions and income levels, consisting of the following image categories: Drinking Water, Drying Clothes, Front Doors, Hand Washing, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes. The dataset captures a mix of actions and static elements, such as home environments, to reflect different aspects of daily life across cultures and income levels.
@@
Prior to obtaining all three annotation sets, the images were selected based on whether ML Objects with at least 50% confidence and ML Captions could be generated. Duplicates were removed, and down-sampling was performed to ensure the images were balanced across regions and income groups. Human annotations were processed to address inconsistencies by removing noise/punctuation, converting to lower case, auto-correcting misspellings, and segmenting concatenated strings. The number of images in the dataset decreased at each step, starting at 3,041 images at the start, reducing to 2,797 after generating the ML Objects and Captions, then to 1,888 after removing duplicates and down-sampling, and finally to 1,886 unique samples after crowdsourcing. In total, 1,155 samples were discarded from the original dataset.
We fine-tuned the pre-trained Sentence Transformer model, “all-MiniLM-L12-v2” (Reimers 2021), using code adapted from Hugging Face’s GitHub repository (Espejel 2022). The model was fine-tuned over 5 epochs with a batch size of 16 to better tailor its representations to our data, rather than using it off the shelf. Fine-tuning on the three annotation sets allowed the model to better capture their characteristics. The resulting sentence embeddings are 384-dimensional, consistent with the original pre-trained model.
Experiment 1: Investigating Annotation Similarity (RQ1)
@@
Experiment 2: Exploring How Annotations Impact Performance and Bias in Predictive Models (RQ2)
This section emphasises the diversity of model prediction outcomes over overall performance. These experiments evaluate the distribution of prediction errors across different regions and categories, allowing us to identify potential discrimination, e.g., higher prediction errors for certain subsets of the population as compared to the error on others. To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
@@
Figure 3 illustrates a t-SNE visualisation of the vector embeddings of the three annotation categories for each data sample. The results show how, despite having a fine-tuned embedding model, the embeddings form into two distinct clusters, with one for ML Captions and Human labels and another for ML Objects. This evidence implies that there exist differences between the different types of annotations we collected in the study. The clusters of embeddings are due to differences in the annotations. Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images. On the other hand, the ML-generated object labels are primarily object-centric. The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification. Moreover, the ML-generated captions would be more descriptive than the object labels and although they may not be as specific as the human annotations, they are more likely to contain actions and objects, which may also be present in Human Labels, hence the clustering of the ML-generated captions and the Human Labels. Additionally, there are apparent syntactic differences between the annotations - the ML Objects consists of lists of objects per image; the ML Captions consist of a phrase per image; while the Human Labels are a list of words and phrases. It is also important to note that a limitation that could have affected this categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information. This causes the sentence structures in the annotations to remain unchanged, resulting in more pronounced categorisations. Considering these differences, the clusters in the visualisation are expected.
@@
Since t-SNE is only a two-dimensional representation of the data, it is useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset. This is shown in Table 1. These values give a more accurate and quantitative comparison of the annotation similarities. It can be seen that the ML Captions and Human Labels are the most similar pair of annotations (0.69), followed by ML Objects and ML Captions (0.50), and ML Objects and Human Labels (0.49). It is important to note that when comparing ML Objects to any of the other annotations, the similarity scores are relatively low. To further investigate length statistics per image, Human Labels averaged 7.18 words per worker, ML Captions averaged 8.62 words per caption, and ML Objects averaged 4.36 object labels per inference run. The length statistic for ML Objects is determined by the number of object labels rather than the number of words because each predicted label corresponds to a single object.
@@
To further explore the relationships of these similarity scores, a plot of pairwise relationships is presented in Figure 4. The plot illustrates how similarity scores between two annotation categories correlate to another pair’s scores for a given image. The plot is colour-coded by region to allow further analysis of pairwise similarity scores by geographical location. The analysis reveals that most regions show no particular trends or patterns, with one exception: images from Europe tend to cluster in the higher similarity range (top right area) of the plots, indicating stronger agreement between all annotation types for European images. Moreover, when looking at the subplots for ML Captions-Human Labels pairs, i.e., subplots on the third column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers. This aligns with the previous observation that ML Captions and Human Labels have the highest average similarity score (0.69). It can also be seen that all three pairs show the same amount of spread of scores across all subplots, which correspond to the standard deviation values of the mean similarity scores between each pair (0.12).
@@
Moreover, the fact that the ML Objects annotations have a relatively low similarity score with the other annotations indicates that the ML object detection model may be detecting objects that do not appear in the corresponding ML Captions or Human Labels, or failing to identify objects that do. This reasoning is consistent with the findings of Zhang et al. (2021), which suggest that deeper neural networks, such as ResNet - the architecture of our object detection model - may struggle more with difficult images. Perhaps there is obstruction or lack of objects in the image, leading to no or incorrect objects being detected in these outliers, resulting in the low similarity scores when compared to their respective ML Captions and Human Labels. The ML captioning model as well as human annotators have the advantage in that they are able to capture non-object elements, such as actions, e.g., “wash”, or environmental descriptions, e.g., “dirty”.
@@
The variation in optimal models across the image categories is understandable. Images from each of the categories may contain varying objects and/or actions that are associated with certain geographical regions, e.g., for Hand Washing, images of buckets may appear more frequently in samples from Africa, whereas images of sinks may appear more in samples from Europe. This explains why certain annotations are better than others at capturing relationships between the objects or actions and the regions. The best-performing models for five out of eight categories (Drying Clothes, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes) are the ones with a combination of at least two annotation sets. This may be due to how the images from these categories may have discriminative features that cannot be captured with just one annotation set.
Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the ML Captions Model (overall best-performing) and the Human Labels Model (overall worst-performing) at $p$-value $<$ $1.15\times 10^{-7}$. This statistically significant value suggests that the ML Captions Model’s better performance is statistically reliable.
@@
It is apparent that, unlike the classification model, the regression model with ML Objects and ML Captions performed the best overall, indicated by its lowest RMSE value (1817.49). The best-performing models differ across all regions. The model containing all annotations was the most effective for African images (561.42), similar to the classification model, while the Asian images achieved the best results with the Human Labels Model (1408.57). Europe saw optimal performance from the ML Objects and ML Captions Model (2777.27), and the Americas was best handled by the ML Captions and Human Labels Model (1578.29). Moreover, it is worth highlighting that Asia is the only region where a stand-alone annotation set achieved the best performance. Looking at the performance of the models by category, the ML Objects-ML Captions Model is the most consistent and achieves the best performance across half of the categories, namely Drinking Water, Drying Clothes, Kitchens, and Washing Clothes. The ML Objects-Human Labels Model was the most effective for Front Doors. The Handwashing category was best handled by the ML Captions Model, while the Human Labels Model achieved the best results for Living Rooms and Places for Dinner.
@@
The observation that the ML Objects and ML Captions Model achieves highest performance aligns with the claim made by Zhang et al. (2021) on machines being better at selecting distinguishing features in the context of image segmentation. In this case of income regression, the superior performance of the ML-based annotations could be attributed to certain objects in the images, as well as descriptive terms indicating the state of the environment (e.g., clean, dirty, old), being associated with certain income levels, therefore improving the model’s ability to predict income. Further experiments would need to be conducted to explore this. Moreover, similar to the classification model, the all-annotations model worked best for African images, likely driven by the need for the annotations to encompass various features, including objects and actions. The images from Asia and the Americas were best handled by the Human Labels Model and the ML Captions-Human Labels Model, respectively. This shows how images from both regions have features indicative of income that are most likely not object-based, rather based on actions or nuanced descriptions of the environment. The fact that Asia required only Human Labels for its optimal performance suggests that the discriminative features for income still require full human intervention, and that although the ML annotations can still capture them, they do not perform as well, even when combined with the Human Labels. In contrast, the European images achieved best performance with the ML Objects-ML Captions Model. This suggests that the combined ML-based annotations are better at capturing the defining features and relationships to predict income for European images. This suggests potential bias in the ML annotations, possible due to Western-centric training data. Alternatively, human annotators may have found better discriminative features for income prediction in images from Asia.
@@
The best regression models for the action-based image categories, e.g., Drying Clothes, Hand Washing, are ones with ML-based annotations, indicating how the ML annotations are better at finding the defining features to predict income. It may be due to how the images contain more easily identifiable actions and objects that are indicative of income. On the other hand, most image categories that are not action-based, such as Living Rooms and Places for Dinner, are best handled by models containing human annotations. This suggests that humans are better at identifying key features for income prediction when said features are likely to be based on nuanced descriptions.
@@
Figure 5 illustrates a scatter plot of the income regression model trained using the ML Objects and ML Captions (best overall) on the test dataset. Each data point represents an individual image sample from the dataset, which corresponds to a household. The points include associated attributes, such as region (continent), country, actual income, and predicted income. They are also colour-coded by continent. The blue dotted line represents the line of best fit and the red solid line represents the identity line. The x-axis represents the ground truth of the income, which has been plotted in log scale for clearer visualisation, whereas the y-axis represents the predicted income, plotted in its original value.
The plot reveals that the incomes of most samples were overestimated, with most of the Africa samples clustering towards the lower income area in the x-axis. On the other hand, the three other regions are more balanced, although there is a great portion of underestimated samples, most of which are from Europe.
Although the data was sampled evenly across continents and balanced across income groups within each continent, the ground truth income levels of most non-African samples are higher than the African samples’. This is evident in the scatter plot, where non-African samples are predominantly positioned to the right of African points. This means that the overall average income across all samples is higher than most African samples, which may attribute to their overestimation. The low variance in the predicted income of African samples may be due to the features in these samples, captured in the ML Objects and ML Captions, being simpler to capture and understand during model training. This simplicity enables the model to be more consistent in its predictions. On the other hand, the key features found in the annotations of the European samples may be more complex, resulting in more variability in predictions, as shown in the vertical spread of the European samples in the scatter plot, as well as the greater variance in the RMSE table.